NASA’s Ames research center in Mountain View, California has a centrifuge built to simulate high-g environments. It spins horizontally in uniform circular motion, with the payload (possibly a human!) in a box at the end of a 29-foot arm. Its maximum rotation rate is 50 revolutions per minute (rpm). (a) What is the centripetal acceleration experienced by an object in the payload box when it is spinning at its maximum rate? (b) How many gs is this acceleration, and could a human survive it (take the maximum survivable g-force to be 8g). (c) What would the rotation rate be to produce this maximum g-force of 8g?

Answer is: a) 240 m/s2; (b) 25g; (c) 29 rev/min
Answer #1

a] Acceleration a = w^2r

= [50*2pi rad/60s]^2*29 foot*0.3048 m/foot as 1foot = 0.3048 m/foot

= 242 m/s^2

b] number of g = 242/9.8 = 24.7g

No human can't survive as it is more than 8g

c] again acceleration a = w^2r

8g = w^2r

8*9.8 = w^2*29*0.3048

w = sqrt((8*9.8)/(29*0.3048))

= 2.98 rad/s

= 2.98 *60/2pi rev/min

= 28.5 rev/min